OpenStudy (anonymous):

please help with polar coordinates. Describe the given region in polar coordinates. The region enclosed by the semicircle x^2 + y^2 = 2y, y >= 0 The solution ends up being r=2sin(theta), 0<(theta)

OpenStudy (fifciol):

first equation of circle x^2-2y +1-1 +y^2=0 (x-1)^2+y^2=1 do you understand this ?

OpenStudy (anonymous):

you moved the y^2 over? where did the 1 and -1 come from

OpenStudy (anonymous):

*2y not y^2

OpenStudy (fifciol):

I added 1 and -1 so I effectively added 0 which doesn't change the expression but it helps to get a nice (y-1)^2 remember: \[y^2-2y+1=(y-1)^2\] there should be in my first post \[(y-1)^2+x^2=1\] not x-1

OpenStudy (anonymous):

I thought you would substitute x=rcostheta and y=rsintheta into the equation x^2+y^2 = 2y?

OpenStudy (fifciol):

No, we're supposed to draw it first to determine the region. Now it's easy to do. We have a circle with center(0,1) and radius r do you agree?

OpenStudy (fifciol):

radius 1 *

OpenStudy (anonymous):

yes

OpenStudy (fifciol):

you have to describe every point on that circumference using r and theta

OpenStudy (fifciol):

sorry theta is wrong i'll type this out

OpenStudy (anonymous):

ok

OpenStudy (fifciol):



OpenStudy (fifciol):

you see now the radius is a function of theta .can you describe the radius in terms of theta?

OpenStudy (fifciol):